OpenStudy (anonymous):

Find all solutions of x^4-5x^3+x-5=0

OpenStudy (anonymous):

facctor by grouping

OpenStudy (anonymous):

take out x^3 to made factor from first 2 term

OpenStudy (anonymous):

Here is what I get x^3(x-5)+(x-5) Then I'm stuck

OpenStudy (anonymous):

now take x-5 out

OpenStudy (anonymous):

so u would get x^3+1(x-5)=0 now put them seprately eaual to zero

OpenStudy (anonymous):

so u get that

OpenStudy (anonymous):

Awesome, I see it now thanks for the help,

OpenStudy (anonymous):

would u have more question?

OpenStudy (anonymous):

So my final answers would be x=5 and x=-1

OpenStudy (anonymous):

ya that's it

OpenStudy (anonymous):

Can you double check this one for me? |7x+6|=3 I get x=-3/7 or x=-9/7 Need to find all solutions

OpenStudy (anonymous):

ya u are right?

OpenStudy (anonymous):

if u need more about then i have a pdf for theory of |....| function .

OpenStudy (anonymous):

Thank you again, I think I am good for now,